# StyleMaster: Towards Flexible Stylized Image Generation with Diffusion Models

### v1

@@ Abstract
Stylized Text-to-Image Generation (STIG) aims to generate images based on text prompts and style reference images. We in this paper propose a novel framework dubbed StyleMaster  for this task by leveraging pretrained Stable Diffusion (SD), which addresses previous problems such as insufficient style and inconsistent semantics. The enhancement lies in two novel modules, namely multi-source style embedder and dynamic attention adapter. In order to provide SD with better style embeddings, we propose the multi-source style embedder, which considers both global and local level visual information along with textual information, which provide both complementary style-related and semantic-related knowledge. Additionally, aiming for better balance between the adapter capacity and semantic control, the proposed dynamic attention adapter is applied to the diffusion UNet in which adaptation weights are dynamically calculated based on the style embeddings. Two objective functions are introduced to optimize the model together with denoising loss, which can further enhance semantic and style consistency. Extensive experiments demonstrate the superiority of StyleMaster  over existing methods, rendering images with variable target styles while successfully maintaining the semantic information from the text prompts.
@@ 1 Introduction
Among all image generation problems, Stylized Image Generation (SIG), in which images with target styles are required to be generated, possesses significant scholarly merit. SIG methods can be smoothly applied to extensive regions such as artistic creation and movie editing, well facilitating real-life cases. Recently, with the surge in diffusion-based generative models [9, 17], the research focus has been transferred from traditional style transfer to the Stylized Text-to-Image Generation (STIG). In this task, one or several style reference images are given as condition, and various images are generated based on the style condition along with additional information such as text prompts. This task, while being more challenging due to mixed input condition, is more flexible and applicable, thus being of higher practical value.
Figure 1: Left: With our proposed StyleMaster , the diffusion-based stylized image generation process can avoid issues such as inconsistent semantics and insufficient style, leading to better generation quality. Right: Our proposed method can generated consistent and delicate stylized images with different styles. For simplicity we only present the generated stylized images here. For the style reference images please refer to the supplementary material.
As a typical STIG method, StyleAdapter [21] is proposed based upon pretrained Stable Diffusion (SD). The reference images are processed with a style embedding module and then injected into the diffusion UNet through extra cross-attention modules, which guide the model to denoise the features according to style information, resulting in stylized images. While such a framework shows promising results for handling specific styles and text prompts, we find that StyleAdapter still suffers from problems such as (1) insufficient style, in which the generated images do not fully share the same style with reference images, for complex styles, and (2) inconsistent semantics, in which semantic objects are leaked from reference images, thus affecting the semantic consistency between the generated images and the text prompts. These problems can be attributed to the design of StyleAdapter. Specifically, StyleAdapter adopts CLIP [16] as source of style information. Such patch-level representation focuses more on local patterns, while ignoring global patterns such as which are also important style descriptors. On the other hand, the adaptation modules in StyleAdapter are applied to all attention layers. In this way, each aggregation with text prompts would be biased by the semantic conuterpart contained in the style embedding, thus resulting in inconsistent semantic.
To this end, we in this paper propose a novel framework for STIG called StyleMaster  with advanced extraction and injection of style information. Our method inherits the basic pipeline of StyleAdapter, while adopting a specific design for style embedding extraction and injection respectively. For style embedding extraction, we propose a multi-source structure to compensate for the patch-level CLIP information. Specifically, we not only use the commonly-used CLIP-based patch embedding to represent image information, but also adopt global-level VGG descriptor, which is engaged in the attention process via adaptive scaling and shifting. Apart from that, we make use of the semantic knowledge of reference images contained in their image captions. This highly abstract semantic information is eliminated from style embedding through the proposed negative embedding branch. By actively merging these different latent spaces, the processed style embedding can thus contain more comprehensive knowledge about the target styles, while avoiding memorizing the semantic content. For style embedding injection, since fully adapting UNet features based on style embeddings can distort the semantic information, a simple solution would be shrinking adapter attachment only to part of UNet, e.g. its upsampling layers. However, direct shrinkage can hurt the capacity of adapters, which in turn worsens the insufficient style problem. To solve this and enhance the module capacity for thorough stylization, we propose the dynamic attention adapter, in which dynamic weights are generated from the style embeddings and then used to adapt both the self-attention and cross-attention layers in the diffusion UNet.
@@
Moreover, we further enhance the model with objectives other than the commonly-used noise prediction loss for diffusion models. We introduce a gram consistency loss, in which we augment the reference images with two sets of transformations, resulting in one set with original style and another set with a disrupted style. Then gram matrices of estimated denoised result and these transformed reference images are calculated as their style-aware statistics. By calculating a triplet loss among them, the model is encourage to generate images with more robust and consistent styles when processing different reference images. Additionally, a semantic disentangle loss is utilized to alleviate the inconsistent semantics problem by contrasting the style embeddings against reference text embeddings, while keeping them similar to reference image embeddings.
@@
To show the effectiveness of our proposed method, we conduct extensive experiments among various styles containing both one-shot and mult-shot settings. We show that our method can significantly outperform baseline methods including StyleAdapter, generating correct styles and avoiding inconsistent semantics. In summary, the contributions of this work are as follows:
  • We propose a novel method for reference-based stylized image generation named StyleMaster, boosting Stylized Text-to-Image Generation (STIG) with a multi-source style embedder and a dynamic attention adapter, greatly solving problems including insufficient style and inconsistent semantics.
  • Two objective functions are introduced to enhance our model, including a gram stylization loss and a semantic disentangle loss, which encourage the model to decouple the style related information from reference images with other style unrelated information.
  • Extensive experiment results show that our method consistently work well in both text-to-image and image-to-image tasks with different reference images, showing the practical value of such a method.
2 Related work
@@
Text-to-image diffusion models. Diffusion models have been proven to be a powerful family of generative models. DDPM [9] originated to propose the framework by modeling the mapping between Gaussian distribution and image distribution with the forward diffusion and inverse denoising process. Based on that Latent Diffusion Model (LDM) [17] largely improved the practical usage by leveraging diffusion model to latent space instead of pixel space, which leads to commonly-known text-to-image diffusion models such as Stable Diffusion (SD), Midjourney and DALLE-3 [1]. Other works focus on improve the diffusion model structure. For example, DiT [14], MDT [6] and PIXART-$\alpha$ [3] utilize the transformer instead of UNet structure, which can be better scaled to larger model size. Blattmann et. al. [2] and Zhang et. al. [25] leverage ideas of Retrieval Augmented Generation (RAG) to generate images based on other retrived images which provide extra knowledge. Yang et. al. [22] propose to leverage the LLMs for planning the text-to-image problems. Our work is built on pretrained SD models, in which the attention mechanism merges conditional information from text prompts to images. Different from the previous works, we focus on designing extra attention adaptation so that the knowledge contained in the style reference images can be smoothly embedded into the denoising process, leading to stylized images.
@@
Stylized image generation. Among all conditional image generation tasks, stylized image generation has long been a highlighted one. Most previous works focus on style transfer, i.e.transfer the style of a content image given another style refernce image. For example, Gatys et. al. [7] solved this problem by optimizing the style-related statistics. Li et. al. [12] equipped this method with explanability. MicroAST [20] proposed to speed up such framework by abandoning the complex visual encoder and utilizing a dual-modulation strategy. Yang et. al. [23] leverages diffusion models, in which the style-aware guidance is used to generate the wanted style. InST [26] realized style transfer by inverting the content image to noise and then re-generate it with the condition control of style images. Apart from these style transfer methods, StyleAdapter [21] proposes a new framework which can generate images directly from style reference images and text prompts without content images. Our work mainly follows StyleAdapter to present a generalized stylization method. Different from StyleAdapter, we analyze the role of style reference images and text prompts in the generation process. Based on that, we propose a novel module to extract more representative style embeddings, which are then injected into noise space with our proposed dynamic adapter.
@@ 3 Preliminary: Stable Diffusion
Diffusion models are probabilistic models that learn a data distribution $p_{\theta}\left(\mathbf{x}_{0}\right)$ by progressively denoising a standard Gaussian distribution, which can be computed by iteratively adding Gaussian noise to the clean data $\mathbf{x}_{0}$. Stable Diffusion (SD) extends such a model to text-to-image based on text prompt $p$. With pre-trained VQ-VAE [19] containing encoder $\mathcal{E}$ and decoder $\mathcal{D}$, SD allows the model to focus more on the semantic information of data and improves efficiency. A diffusion UNet is used to predict the noise, in which attention mechanism is adopted. Specifically, for the $l$-th layer, self-attention is first used to interact among spatial features: $z^{l}=Attention(W_{Q}^{l}\cdot z^{l},W_{K}^{l}\cdot z^{l},W_{V}^{l}\cdot z^{l})$, where $Attention$ denotes the attention operator, $z^{l}$ denotes latent embeddings of the $l$-th layer, $W_{Q},W_{K},W_{V}$ denotes the projection layers of self-attention. After that the cross-attention is utilized to merge condition information such as text prompt: $\hat{z}^{l}=Attention(\hat{W}_{Q_{t}}^{l}\cdot z^{l},\hat{W}_{K_{t}}^{l}\cdot z_{text},\hat{W}_{V_{t}}^{l}\cdot z_{text})$, where $z_{text}$ denotes text prompt embedding, $\hat{W}_{Q},\hat{W}_{K},\hat{W}_{V}$ denotes the projection layers of cross-attention. The training objective of SD is as follows:
@@ 4 Methodology
Figure 2: Overview of our proposed StyleMaster. Concretely, the multi-source style embedder (Sec. 4.1) aggregates style patterns contained in each reference image simultaneously from global and local levels. Style embeddings are then engaged in the denoising procedure through the proposed dynamic attention adaptation (Sec. 4.2), which guides both the attentions in diffusion UNet to properly merge style and semantic information from different sources. During training augmented input images are used as style reference images, through which objectives as described in Sec. 4.3 are used to optimize the model. During inference the style reference images are achieves with manual assignment instead of using augmentation of a specific image.
@@
We focus on reference-based Stylized Text-to-Image Generation (STIG) in this paper. Formally, a reference style image set $\mathcal{I}_{s}=\{\mathbf{I}_{style}^{i}\}_{i=1}^{N_{s}}$, where $N_{s}$ denotes the number of reference images, together with text prompt $p$ are given as condition information. $N_{s}$ can be variable among different trials to describe different style concepts. The model is required to generate image $\mathbf{I}$ that shares the same style pattern with $\mathcal{I}_{s}$ and same semantic meaning with text prompt $p$. To solve this task we present a novel framework named StyleMaster  based on SD, as shown in Fig. 2, which will be introduced in this section.
Figure 3: Left: multi-source style embedder. Right: dynamic attention adapter.
4.1 Multi-source style embedder
Given the successful application of text-to-image based on SD, an ideal reference-based stylization should necessarily depend on a representative style embeddings. StyleAdapter [21] adopts CLIP visual encoder [16] to extract the patch-level features from style reference images, which are then processed with cross-attentions with learnable style tokens. However, StyleAdapter still suffers from problems like insufficient style and inconsistent semantics, i.e.the semantic knowledge in reference images are generated during the denoising procedure. The former problem can be attributed to that CLIP-based patch embeddings concentrate more on local patterns. In this way, the style embeddings fail to hold the global style patterns, leading to insufficient style. Meanwhile, since CLIP image embeddings are aligned with their captions, they inevitably contain information about contents in the images. Engaging such information in style embeddings can lead to inconsistent semantics.
@@ 4.1 Multi-source style embedder
To solve these problems, we propose to leverage knowledge from multiple sources to complement the style embedding. Specifically, given a style image set $\mathcal{I}_{s}$, we extract three different kinds of features with pretrained models: (1) local style-aware features: Following StyleAdapter, we utilize CLIP to encode each style image $\mathbf{I}_{style}^{i}$ into latent patch tokens $z_{CLIP}^{i}$, which are then concatenated together along the token dimension, resulted as $z_{CLIP}$. (2) global style-aware features: Inspired by previous style transfer methods [7], we adopt the gram matrix to represent global-level style information. Specifically, $\mathbf{I}_{style}^{i}$ is processed with VGG-19 [18] to get the relu3_1 feature $z_{vgg}^{i}$, using which the gram matrix can be calculated as $z_{gram}^{i}={z_{vgg}^{i}}^{T}\cdot z_{vgg}^{i}$. Then gram matrices for different style images are averaged to get the final representation $z_{gram}$. (3) semantic-aware features: We adopt CLIP text encoder to extract the text embedding $z_{cap}^{i}$ of captions of $\mathbf{I}_{style}^{i}$, which is then concatenated into $z_{cap}$. During training, the captions are provided in the training set, while during inference, we adopt BLIP [11] to first annotate the style images.
@@
Generally, each token in $z_{CLIP}$ contains local-level style features, while $z_{gram}$ describes more abstract style information. On the other hand, $z_{cap}$ contains the semantic information of $\mathcal{I}_{s}$, which should be eliminated in the final style embedding to avoid inconsistent semantics. To properly make use of these embeddings, we propose a novel dual-branch structure as shown in Fig. 3(a). Concretely, several learnable style tokens $z_{s}$ are first attached to $z_{CLIP}$, with their replication $z^{N}_{s}$ denoted as negative semantic tokens attached to $z_{cap}$:
@@
where $\cdot\|_{t}\cdot$ denotes concatenation along the token dimension, $\delta_{t}$ is the same time embedding of denoising timestep as used in diffusion UNet. $\hat{z}_{caption}$ is then individually processed with several transformer layers to aggregate the information between style tokens and text embedding. For $\hat{z}_{CLIP}$, we adopt a modified version of transformer layer. Specifically, $z_{gram}$ is first projected to scaling and shift coefficients. These coefficients are applied to the self-attention procedure in the same way as adaLN [15]. Before processing the attention result with FFN, the style token part in $\hat{z}_{caption}$ is subtracted from the counterpart in $\hat{z}_{CLIP}$.
Such design enjoys three major merits. First, the module is timestep-aware. Since the style tokens are merged with time embedding, different denoising steps can thus model different information. Second, the global information in $z_{gram}$ can guide the model to better concentrate on style-related knowledge, thus avoiding those irrelevant but repetitive local patterns shared among style images. Third, the negative semantic tokens generally contain more abstract content information rather than style. Consequently, subtracting it from $\hat{z}_{CLIP}$ can help alleviate the problem of inconsistent semantics. While some captions may describe the style of images, the model can learn to maintain these knowledge during training thanks to the training strategy described in Sec. 4.3. By using the proposed module, we can learn more representative and generalizable style embeddings which can better facilitate the stylization process described as follow.

### v2

# ArtWeaver: Advanced Dynamic Style Integration via Diffusion Model

@@ Abstract
Stylized Text-to-Image Generation (STIG) aims to generate images from text prompts and style reference images. In this paper, we present ArtWeaver, a novel framework that leverages pretrained Stable Diffusion (SD) to address challenges such as misinterpreted styles and inconsistent semantics. Our approach introduces two innovative modules: the mixed style descriptor and the dynamic attention adapter. The mixed style descriptor enhances SD by combining content-aware and frequency-disentangled embeddings from CLIP with additional sources that capture global statistics and textual information, thus providing a richer blend of style-related and semantic-related knowledge. To achieve a better balance between adapter capacity and semantic control, the dynamic attention adapter is integrated into the diffusion UNet, dynamically calculating adaptation weights based on the style descriptors. Additionally, we introduce two objective functions to optimize the model alongside the denoising loss, further enhancing semantic and style consistency. Extensive experiments demonstrate the superiority of ArtWeaver over existing methods, producing images with diverse target styles while maintaining the semantic integrity of the text prompts.
Figure 1: Results generated by StyleAdapter and our proposed ArtWeaver. Our method can better stylized the images based on the reference images. Compared with StyleAdapter, our method can generate more consistent and delicate stylized images with different styles. For simplicity we only present the generated images here. For simplicity only one reference image for each style is shown. For the full reference images please refer to the supplementary material.
@@ 1 Introduction
Stylized Image Generation (SIG), which involves generating images with specific artistic styles, has significant academic and practical implications, particularly in fields such as art and film. The emergence of diffusion-based generative models [10, 19] has shifted the focus from traditional style transfer techniques to Stylized Text-to-Image Generation (STIG). In STIG, one or more style reference images are used as conditions to generate various images, incorporating additional information such as text prompts. Despite the complexity of handling mixed input conditions, STIG methods offer enhanced flexibility, making them highly applicable to real-world scenarios.
Recent advancements in Stable Diffusion (SD) based STIG methods, such as StyleAdapter [25] and InstantStyle [24], have demonstrated promising few-shot stylization capabilities. These methods extract style descriptors from reference images using a style embedding module, which are then injected into the diffusion UNet via cross-attention modules to guide the stylization process during denoising. This framework has been further extended in works like ArtAdapter [3] and DEADiff [17]. However, these methods still encounter significant challenges: (1) misinterpreted style, where the generated images do not fully capture the intricate styles of the reference images, and (2) inconsistent semantics, where elements from the reference images improperly influence the output, leading to misalignment with the text prompts.
These challenges primarily arise from the design of the style embedding extractor. Traditional methods often rely on a single-source style descriptor, using either pretrained models like CLIP [18] or custom networks. However, reference images typically contain multi-level information, including local textures and global color schemes, along with rich semantic content which can bias the styles. Moreover, single-source style descriptors mix both low-frequency and high-frequency content, even though different denoising steps need different types of information. Consequently, such an approach can limit the style descriptor’s ability to represent styles accurately, leading to less effective performance. Additionally, injecting style embeddings into all cross-attention layers of the diffusion UNet can bias the model’s integration with text prompts, resulting in semantic inconsistencies.
In this paper, we introduce ArtWeaver, a novel STIG framework that features advanced techniques for extracting and injecting style information. Our method builds upon the foundational pipeline of StyleAdapter, incorporating specific enhancements to improve style embedding extraction and injection. For style extraction, we propose a Mixed Style Descriptor (MSD). First, we follow StyleAdapter to adopt the CLIP-based patch descriptor, based on which we apply disentanglement according to the frequency domain. In addition, we incorporate a gram-based style descriptor via a Gram matrix [8] and a semantic descriptor derived from reference image captions. These diverse sources of information are integrated through transformer layers with adaptive scaling and shifting, where the semantic descriptor is aligned with and subtracted from the other descriptors. This approach captures a more comprehensive representation of target styles while preventing semantic leakage. For style embedding injection, a naive strategy as adopted by InstantStyle [24] is to limit the attachment of adapters to specific parts of the UNet, such as the upsampling layers, to prevent semantic distortion. However, this direct limitation can reduce the capacity of adapters and exacerbate the issue of misinterpreted style. To address this, we propose a Dynamic Attention Adapter (DAA) that generates sample-specific dynamic weights from the style embeddings. Compared with directly employing more layers for cross-attention modules, DAA leverages current intermediate features to adapt both self-attention and cross-attention layers in the diffusion UNet, allowing for more precise and flexible style adaptation. This ensures that the generated images maintain both the intended style and semantic consistency with the text prompts.
@@
Furthermore, we enhance the model with objectives beyond the standard noise prediction loss commonly used in diffusion models. We introduce a Gram consistency loss, augmenting the reference images with two sets of transformations: one set preserves the original style, while the other adopts a distorted style. We then compute Gram matrices of the estimated denoised results and these transformed reference images as their style-aware statistics. By applying a triplet loss among these matrices, the model is encouraged to generate images with more robust and consistent styles when processing different reference images. Additionally, we utilize a semantic disentanglement loss to mitigate the inconsistent semantics problem by contrasting the style embeddings against reference text embeddings, while ensuring they remain similar to the reference image embeddings.
@@
To demonstrate the effectiveness of our proposed method, we conduct extensive experiments across various styles, including both one-shot and multi-shot settings. Our results show that ArtWeaver significantly outperforms baseline methods like StyleAdapter, generating accurate styles and avoiding inconsistent semantics. In summary, the contributions of this work are as follows:
1) Our ArtWeaver introduce the Mixed Style Descriptor (MSD) which captures comprehensive target style representations while preventing semantic leakage through a negative embedding branch.
2) To enhance style embedding injection, ArtWeaver introduces a dynamic attention adapter that generates weights from style embeddings, enabling precise adaptation of self-attention and cross-attention layers in the diffusion UNet.
3) Our model incorporates objectives beyond standard noise prediction loss, including a novel Gram consistency loss that promotes robust and consistent styles through triplet loss on Gram matrices from transformed reference images. Additionally, a semantic disentanglement loss contrasts style embeddings with reference text embeddings while maintaining similarity to reference image embeddings, addressing inconsistent semantics.
2 Related Works
@@
Text-to-image diffusion models. Diffusion models have been proven to be powerful generative models. DDPM [10] originated to propose the framework by modeling the mapping between Gaussian distribution and image distribution with the forward diffusion and inverse denoising process. Based on that Latent Diffusion Model (LDM) [19] largely improved the practical usage by leveraging diffusion model to latent space instead of pixel space, which leads to commonly-known text-to-image diffusion models such as Stable Diffusion (SD), Midjourney and DALLE-3 [1]. Other works focus on improve the diffusion model structure. For example, DiT [15], MDT [7] and PIXART-$\alpha$ [4] utilize the transformer instead of UNet structure, which can be better scaled to larger model size. [2] and [29] leverage ideas of Retrieval Augmented Generation (RAG) to generate images based on other retrived images which provide extra knowledge. [27] propose to leverage the LLMs for planning the text-to-image problems. Different from the previous works based on SD, we focus on designing extra attention adaptation so that the knowledge contained in the style reference images can be smoothly embedded into the denoising process, leading to stylized images.
@@
Stylized image generation. Among all conditional image generation tasks, stylized image generation has long been a highlighted one. Most previous works focus on style transfer, i.e., transfer the style of a content image given another style reference image. For example, MicroAST [26] proposed to speed up such framework by abandoning the complex visual encoder and utilizing a dual-modulation strategy. InST [30] realized style transfer by inverting the content image to noise and then re-generate it with the condition control of style images. StyleAdapter [25] and ArtAdapter [3] proposed a new framework which can generate images directly from style reference images and text prompts without content images. StyleID [5] adopted training-free approach to inverse both content image and style image into noises than merge them. DEADiff [17] utilizes a pipeline similar to StyleAdapter, but leverages a self-constructed high-quality dataset. Somepalli et. al. [22] proposed a style descriptor with contrastive learning. StyleTokenizer [13] built a new tokenizer to align the latent space of style and textual information. Our work mainly follows StyleAdapter to present a generalized stylization method. Different from StyleAdapter, we analyze the role of style reference images and text prompts in the generation process. Based on that, we propose a novel module to extract more representative style embeddings, which are then injected into noise space with our proposed dynamic adapter.
@@ 3 Preliminary: Stable Diffusion
Diffusion models model the data distribution $p_{\theta}\left(\mathbf{x}_{0}\right)$ of clean data $\mathbf{x}_{0}$ by progressively denoising a standard Gaussian distribution, of which the learning process is instantiated as denoising score matching. Stable Diffusion (SD) extends such a model to text-to-image based on text prompt $p$. With pre-trained VQ-VAE [23] containing encoder $\mathcal{E}$ and decoder $\mathcal{D}$, SD allows the model to focus more on the semantic information of data and improves efficiency. A diffusion UNet is used to predict the noise, in which attention mechanism is adopted. Specifically, for the $l$-th layer, self-attention is first used to interact among spatial features: $z^{l}=Attention(W_{Q}^{l}z^{l},W_{K}^{l}z^{l},W_{V}^{l}z^{l})$, where $Attention$ denotes the attention operator, $z^{l}$ denotes latent embeddings of the $l$-th layer, $W_{Q},W_{K},W_{V}$ denotes the projection layers of self-attention. After that the cross-attention is utilized to merge condition information such as text prompt: $\hat{z}^{l}=Attn(\hat{W}_{Q_{t}}^{l}z^{l},\hat{W}_{K_{t}}^{l}z_{text},\hat{W}_{V_{t}}^{l}z_{text})$, where $z_{text}$ denotes text prompt embedding, $\hat{W}_{Q},\hat{W}_{K},\hat{W}_{V}$ denotes the projection layers of cross-attention. The training objective of SD is as follows:
@@ 4 Methodology
Figure 2: Overview of our proposed ArtWeaver. Concretely, the Mixed Style Descriptor (MSD) (Sec. 4.1) aggregates style patterns contained in each reference image simultaneously from global and local levels. Style embeddings are then engaged in the denoising procedure through the proposed Dynamic Attention Adaptation (DAA) (Sec. 4.2), which guides both the attentions in diffusion UNet to properly merge style and semantic information from different sources. During training augmented input images are used as style reference images, through which objectives as described in Sec. 4.3 are used to optimize the model. During inference the style reference images are achieved with manual assignment instead of using augmentation of a specific image.
@@
We focus on reference-based Stylized Text-to-Image Generation (STIG) in this paper. Formally, a reference style image set $\mathcal{I}_{s}=\{\mathbf{I}_{style}^{i}\}_{i=1}^{N_{s}}$, where $N_{s}$ denotes the number of reference images, together with text prompt $p$ are given as condition information. $N_{s}$ can be variable among different trials to describe different style concepts. The model is required to generate image $\mathbf{I}$ that shares the same style pattern with $\mathcal{I}_{s}$ and same semantic meaning with text prompt $p$. To solve this task we present a novel framework named ArtWeaver based on SD, as shown in Fig. 2, which will be introduced in this section.
Figure 3: The structure of the Mixed Style Descriptor (MSD)
4.1 Mixed Style Descriptor
Given the successful application of text-to-image generation based on Stable Diffusion (SD), an ideal reference-based stylization should rely on extracting style embeddings that are as representative as the text embeddings produced by representative models like CLIP and T5. However, previous methods still face issues such as misinterpreted styles and inconsistent semantics. These issues arise because the previously adopted style descriptor is limited to both comprehensively represent the style and get rid of the negative effect of semantic information. Moreover, as SD models different kinds of information across various denoising timesteps, providing identical and entangled frequency-related information forces SD to extract the required information independently, increasing its learning burden.
@@ 4.1 Mixed Style Descriptor
To address these problems, we introduce a comprehensive Mixed Style Descriptor (MSD). Specifically, given a style image set $\mathcal{I}_{s}$, we extract three different types of features using pretrained models: (1) CLIP-based style descriptor: Following StyleAdapter, we use CLIP to encode each style image $\mathbf{I}_{style}^{i}$ into latent patch tokens $\tilde{z}_{CLIP}^{i}\in\mathbb{R}^{c\times(wh)}$, where $c$, $h$, and $w$ denote the latent channel, height, and width of CLIP features. We then apply discrete wavelet transform (DWT) to $\tilde{z}_{CLIP}^{i}$, resulting in low-frequency features $\tilde{z}_{CLIP,lf}^{i}$ and high-frequency counterparts $\tilde{z}_{CLIP,hf}^{i}$. All low and high-frequency features from $\mathcal{I}_{s}$ are concatenated along the token dimension, forming $z_{CLIP}$. (2) Gram-based style descriptor: Inspired by Neural Style Transfer (NST) [8], we use the Gram matrix to complement the style information. Specifically, following NST, $\mathbf{I}_{style}^{i}$ is processed with pretrained VGG-19 [20] to obtain the relu3_1 feature $z_{vgg}^{i}\in\mathbb{R}^{c^{\prime}\times h^{\prime}w^{\prime}}$, where $c^{\prime}$, $h^{\prime}$, and $w^{\prime}$ denote the latent channel, height, and width of the VGG feature map. The Gram matrix is then calculated as $z_{gram}^{i}=z_{vgg}^{i}{z_{vgg}^{i}}^{T}$, which is flattened to a vector with dimension $\mathbb{R}^{c^{2}}$. The Gram matrices for different reference images are averaged to obtain the final representation $z_{gram}$. (3) Semantic descriptor: We use the CLIP text encoder to extract the text embedding $z_{cap}^{i}$ of the captions of $\mathbf{I}_{style}^{i}$, which are then concatenated into $z_{cap}$. During training, the captions are provided in the training set. During inference, we use BLIP [12] to annotate the caption of style reference images, which can be replaced with other advanced image caption models for future works.
@@
Generally, each token in $z_{CLIP}$ contains style features related with image content and disentangled in the frequency domain, while $z_{gram}$ focuses less on image content but more on global statistics related with the style of interest. On the other hand, $z_{cap}$ contains the semantic information of $\mathcal{I}_{s}$, which should be eliminated in the final style descriptor to avoid inconsistent semantics. To properly make use of these descriptors, we propose a novel dual-branch structure as shown in Fig. 3. Concretely, several learnable style tokens $z_{s}$ are first attached to $z_{CLIP}$, with their replication $z^{N}_{s}$ denoted as negative semantic tokens attached to $z_{cap}$:
@@
where $\|_{t}$ denotes concatenation along the token dimension, $\delta_{t}$ is the same time embedding of the denoising timestep as used in diffusion UNet. $\hat{z}_{caption}$ is then individually processed with several transformer layers to aggregate the information between style tokens and text embedding. For $\hat{z}_{CLIP}$, we adopt a modified version of transformer layer. Specifically, $z_{gram}$ is first projected to scaling and shift coefficients. These coefficients are applied to the self-attention procedure the same as adaLN [16]. Before processing the attention result with FFN, the style token part in $\hat{z}_{caption}$ is projected with MLP to align with the image latent space and subtracted from its counterpart in $\hat{z}_{CLIP}$.
Our design offers three major advantages. First, the module is timestep-aware. Since the frequency-aware style tokens are merged with the time embedding, different denoising steps can model different information. Second, $z_{gram}$ can well complement the information provided by $z_{CLIP}$ to better focus on style-related knowledge, thus avoiding irrelevant but repetitive local contents shared among style images. Third, the negative semantic tokens generally contain more abstract content information rather than style. Consequently, subtracting them from $\hat{z}_{CLIP}$ can help alleviate the problem of inconsistent semantics. While some captions may describe the style of images, the model can learn to maintain this knowledge during training thanks to the training strategy described in Sec. 4.3. By using the proposed module, we can learn more representative and generalizable style embeddings, which can better facilitate the stylization process described as follows.
Figure 4: The structure of the Dynamic Attention Adapter (DAA)